OpenStudy (anonymous):

write the solution in interval notation x<5 and x ≥ -2

OpenStudy (anonymous):

for what basic set?

OpenStudy (anonymous):

x<5 and x ≥ -2

OpenStudy (anonymous):

no i meant the set... real numbers? natural numbers?

OpenStudy (anonymous):

this is all i have to work with and a graph

OpenStudy (anonymous):

\[(\infty,5)\]\[[-2, \infty)\]

jhonyy9 (jhonyy9):

for x<5 will be -infinity till 5 but five not can being for x>=-2 will be -2 +infinity inclusive -2 too

OpenStudy (anonymous):

oh snap... sry - the first one was supposed to be\[(-\infty, 5)\]sorry about that. Thanks jhonyy9 for pointing it out! :)